OpenStudy (anonymous):

Describe an infinite geometric series with a beginning value of 2 that converges to 10. what are the first 4 terms of the series?

OpenStudy (anonymous):

Please Help

OpenStudy (welshfella):

sum to infinty is a1/(1-r) where a1 = ffirst term and r is the common difference

OpenStudy (welshfella):

so we have the relation 10 = 2 (1 - r) Solve for r.

OpenStudy (welshfella):

sorry that is 10 = 2 / (1 - r)

OpenStudy (welshfella):

so 10 - 10r = 2 10r = 8 r = 0.8

OpenStudy (welshfella):

so the first 3 terms are 2 , 2*0.8, 2*0.8^2

OpenStudy (anonymous):

ok so i would multiply 2*0.8 and 2*0.8^2 right? and then so on?

OpenStudy (anonymous):

@welshfella